OpenStudy (anonymous):

Solve the systems by addition method. 3x-2y=19 2x-3y=11

OpenStudy (aripotta):

is addition method the elimination method?

OpenStudy (anonymous):

3x-2y=19 <---- -2(3x-2y=19) --->-6x+4y=-38 2x-3y=11 <----- 3(2x-3y=11) ---> 6x-9y = 33 -------------------------------------------------- -5y=-5 --> y=1

OpenStudy (anonymous):

but when i go plug in 1 for y it doesnt come out -38

OpenStudy (anonymous):

im not done to i have to solve for x right!??

OpenStudy (aripotta):

yea, plug in 1 for y in 3x - 2y = 19. then solve for x

OpenStudy (anonymous):

dang (7,1) i need to get some rest!

OpenStudy (anonymous):

@Mandy86 your procedure is perfectly right... x will never be 38...

OpenStudy (anonymous):

thanks guys I know how to do this its just im so tired

OpenStudy (anonymous):

x=7,y=1